Find the absolute maximum and absolute minimum of the function f(x,y)=2x^3+y^4 on the region {(x,y)|x^2+y^2≤4}.

If the absolute max or min occurs at multiple points list them all, separated by commas.

Absolute minimum value: ___

attained at ___.

Absolute maximum value: ___

attained at ___.
1
Expert's answer
2020-10-21T17:25:45-0400

Absolute max and min "f(x,y)=2x^3+y^4" on region "x^2+y^2\leq4"


first order partial derivatives:

"f_x = 6x^2, f_y=4y^3"

"f_x=0, f_y=0\implies x=0, y=0"

critical point (0,0) is inside the disk of radius 2

the function value of this critical point is "f(0,0)=0"


checking the boundaries:

"x^2+y^2=4"

"y^2=4-x^2"

"-2\leq x \leq 2"

"g(x)=2x^3+(4-x^2)^2=2x^3+16-8x^2+x^4=x^4+2x^3-8x^2+16"

"g'(x)=4x^3+6x^2-16x \implies x=\frac{-3\pm\sqrt{73}}{4}" or 0

"x_1=0,x_2\approx-2.886, x_3\approx1.386"

the value of g(x) at the critical points and the endpoints:

"g(-2)=-16, g(2)=16, g(0)=16, g(-2.886)\approx-29.33, g(1.386)\approx9.65"

"x=-2: y^2=4-x^2=4-4=0 \implies y=0"

"x=2: y^2=4-x^2 \implies y=0"

"x=0: y^2=4-0=4 \implies y=\pm2"

"x=-2.886: y^2=4-x^2=-4.33 \implies\nexists y"

"x=1.386: y^2=4-x^2=2.08 \implies y=\pm1.44"


The function values for g(x) correspond to the following function values for f(x,y):

"g(-2)=-16 \implies f(-2,0)=-16"

"g(2)=16 \implies f(2,0)=16"

"g(0)=16 \implies f(0,2)=16, f(0,-2)=16"

"g(1.386)=9.65 \implies f(1.386,1.44)=9.65, f(1.386,-1.44)=9.65"


Comparing these values we finally have the answer:

absolute minimum value is -16 and occurs at (-2,0)

absolute maximum value is 16 and occurs at (2,0), (0,2) and (0,-2)